How could I find the exact value of cos(sin^-1(3/5)?
Answer by stanbon(48564)

How could I find the exact value of cos(sin^-1(3/5)?
---------
sin^-1(3/5) is an angle in the 1st quadrant because sin^-1 has a range of
-90 to 90.
x, y, and r are all positive in the 1st quadrant.
You are given opp/hyp = 3/5; you need to find adjacent.
adj=sqrt(hyp-opp) = sqrt(5^2 - 3^2) = sqrt(16) = 4
Therefore cos(sin^-1(3/5))= 4/5
--------
Cheers,
Stan H.